# Accelerating Natural Gradient Descent for PINNs with Randomized Numerical Linear Algebra

### v2

@@ 4 NyströmNGD: Matrix-Free Natural Gradient Descent with Randomized Nyström Preconditioning
To compute the regularized natural gradient, we approximately solve $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})\mathbf{d}=\nabla L(\boldsymbol{\theta})$ using the preconditioned conjugate gradient (PCG) method in a matrix-free fashion, with the preconditioner constructed via randomized Nyström. This approach leverages the structure of the problem: matrix-vector products with the Gramian $\mathbf{G}(\boldsymbol{\theta})$ can be computed efficiently via automatic differentiation, while the strong spectral decay commonly observed in practice (see Figure 1) permits an accurate low-rank approximation, yielding a preconditioner that is both efficient to apply and effective in accelerating convergence. This section presents the resulting algorithm in full detail.
Choice of the regularization parameter $\mu$
@@
The regularization parameter $\mu>0$ is crucial for numerical stability and for mitigating the slow initial convergence typical of quasi-second-order methods like NGD. Prior works proposed heuristics that adapt $\mu$ dynamically [12, 28] or decrease it with the loss [72], so that early iterations mimic gradient descent and gradually shift to NGD near a local minimum. We instead propose a heuristic based on the condition number of $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})$.
@@
Letting $\lambda_{1}\geq\dots\geq\lambda_{p}$ denote the eigenvalues of the Gramian $\mathbf{G}(\boldsymbol{\theta})$, the spectral condition number of $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})$ is $k=(\lambda_{1}+\mu)/(\lambda_{p}+\mu)\leq 1+\lambda_{1}/\mu$. Hence, choosing $\mu=c\cdot\lambda_{1}$ guarantees $k\leq 1+c^{-1}$. In practice, good results are obtained by setting $c$ proportional to machine epsilon, i.e., $\mu=\gamma\cdot\epsilon_{\mathrm{mach}}\cdot\lambda_{1}$, with $\gamma=p$ as a reasonable default. Observe that although $\lambda_{1}$ is not directly available, it can be estimated by $\hat{\lambda}_{1}$ from randomized Nyström. It can be shown that the expected relative error can be small even for modest values of $\ell$, when $\mathbf{G}(\boldsymbol{\theta})$ exhibits strong spectral decay [51, Theorem 14.1].
Choice of rank parameter $\ell$
@@
The rank parameter $\ell$ controls the computational cost of the preconditioner and should ideally scale with the effective dimension $d_{\mathrm{eff}}(\mu)$ [26, Theorem 1.1]. As $d_{\mathrm{eff}}(\mu)$ is not directly accessible, we estimate $\ell$ adaptively. Following [26, Section 5.4.2], we do it by monitoring the ratio $\hat{\lambda}_{\ell}/\mu$: if it drops below $10$, we double $\ell$ for the next iteration, up to a prescribed maximum $\ell_{\max}$. Otherwise, we reduce $\ell$ to the smallest index satisfying $\hat{\lambda}_{\ell}/\mu<10$, adding a small offset (e.g., $1$ or $2$) to avoid oscillations.
@@
Algorithm 1 summarizes the proposed algorithm. At each iteration, we construct the randomized Nyström preconditioner (lines 5 to 7) using the matrix-free Gramian. We then apply preconditioned conjugate gradient (PCG) to solve the linear system (3) (line 9). The PCG stopping criterion combines a maximum number of iterations maxit and a relative tolerance $\mathtt{rel\_tol}=\min(\kappa,\left\|{\nabla L(\boldsymbol{\theta}_{k})}\right\|_{2})$, inspired by the Steihaug–Toint truncated CG method [18]. This design avoids overspending computation far from critical points, where NGD steps may underperform compared to GD, and ensures accuracy near local minima. Once the search direction $\mathbf{d}_{k}$ is computed, a line search is performed to update the iterate (lines 10 and 11). In practice, we found that a standard backtracking line-search [70] suffices, without requiring logarithmic grid searches as in [54, 38]. Finally, the rank parameter $\ell$ is updated adaptively (lines 12 to 16).
@@
5:     $[\mathbf{U},\hat{\boldsymbol{\Lambda}}]=\mathtt{RandomizedNystrom}(\mathbf{G}_{\mathrm{fun}},\ell)$triangleright Randomized Nyström approximation
6:     $\mu=\gamma\cdot\hat{\lambda}_{1}\cdot\epsilon_{\mathrm{mach}}$triangleright Regularization parameter $\mu$
7:     $\mathbf{P}_{\mathrm{fun}}=\mathbf{v}\mapsto\frac{1}{\hat{\lambda}_{\ell}+\mu}\mathbf{U}(\hat{\boldsymbol{\Lambda}}+\mu\mathbf{I})\mathbf{U}^{\top}\mathbf{v}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\top})\mathbf{v}$triangleright Matrix-free preconditioner
8:     $\mathtt{rel\_tol}=\min(\kappa,\left\|{\nabla L(\boldsymbol{\theta}_{k})}\right\|_{2})$triangleright Relative tolerance for PCG
9:     $\mathbf{d}_{k}=\texttt{PCG}(\mathbf{G}_{\mathrm{fun}}+\mu\cdot\mathrm{Id},\,\nabla L(\boldsymbol{\theta}_{k}),\,\mathtt{rel\_tol},\,\mathtt{maxit},\,\mathbf{P}_{\mathrm{fun}})$triangleright Run PCG
10:     $\alpha_{k}=\texttt{Linesearch}(\boldsymbol{\theta}_{k},-\mathbf{d}_{k})$triangleright Linesearch along $\alpha\mapsto\boldsymbol{\theta}_{k}-\alpha\mathbf{d}_{k}$
@@
12:     if $\hat{\lambda}_{\ell}<10\mu$ then
13:         $\ell=\min(2\ell,\ell_{\max})$triangleright Possibly increase rank parameter $\ell$
@@
15:         $\ell=\min\left(\min\{i:\hat{\lambda}_{i}>10\mu\}+1,\,\ell_{\max}\right)$triangleright Possibly decrease rank parameter $\ell$
@@ 5 Numerical Experiments
We test NyströmNGD on a range of forward and inverse PDE problems with known analytical solutions. Additional resources and implementation details are provided in Appendix E. Additional numerical experiments involving variational formulations of PDEs are presented in Appendix F.
@@
We compare NyströmNGD against Natural Gradient Descent with SVD-based pseudoinverse (legend: NGD), matrix-free NGD solved via conjugate gradient without preconditioning (legend: NGD with CG), and the BFGS algorithm [70]. We run NGD-based methods for $500$ or $1000$ iterations, depending on the computational cost, and BFGS for $6\times 10^{4}$ or $1.2\times 10^{5}$ iterations. For a fair comparison with NyströmNGD, in NGD with CG we use the same relative tolerance as in Algorithm 1 and set the maximum number of iterations to $\texttt{maxit}+\ell_{\max}$. Given the different per-iteration cost of the methods, we compare them in both iteration count and wall-clock time. We do not include first-order optimizers (e.g., gradient descent, Adam) or L-BFGS, as they have been shown to be non-competitive [38, 55]. In all experiments, we employ fully connected neural networks with $\tanh$ activation and two hidden layers of width $32$. Integrals in both the loss and the metric are discretized via Monte Carlo quadrature. All experiments are averaged over 10 runs. In the plots, the solid line denotes the median, and the shaded region spans the first and third quantiles.
@@
For the PINN formulations, we use JAX [15] and JAXopt [11]. The JAXopt BFGS implementation was optimized by us for improved efficiency, as detailed in Appendix D, yielding, toghether with JIT compilation, a 1000×1000$\times$1000 × speedup over the timings reported in [54]. The NGD implementation is taken from the public repository of [72]. All experiments were conducted in double precision on a single NVIDIA RTX 4090 GPU with 24GB of memory.
5.1 Poisson problem in 3D
@@
with exact solution $u(x,y,z)=\sin(\pi x)\sin(\pi y)\sin(\pi z)$ as in [72]. The PINNs least-squares loss and the corresponding metric are
$E(u)=\frac{1}{2}\left\|{\Delta u+f}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}+\frac{1}{2}\left\|{u-g_{D}}\right\|_{\mathrm{L}^{2}(\partial\Omega)}^{2},$
${g}\left({u},\,{v}\right)=\left({\Delta u},\,{\Delta v}\right)_{\mathrm{L}^{2}(\Omega)}+\left({u},\,{v}\right)_{\mathrm{L}^{2}(\partial\Omega)}=\frac{1}{2}\int_{\Omega}\Delta u\,\Delta v\,\mathrm{d}\mathbf{x}+\frac{1}{2}\int_{\partial\Omega}u\,v\,\mathrm{d}\mathbf{s}.$
Further details are provided in Section E.1.
@@
As shown in Figure 2, NyströmNGD yields the best results in both final accuracy and computational efficiency, converging in around $100$ iterations. The Nyström rank parameter $\ell$ increases rapidly during early iterations, consistent with Figure 1, and then stabilizes well below the maximum allowed rank $\ell_{\max}=500$. This confirms the effectiveness of the adaptive rank selection. NGD initially converges faster in iterations but plateaus early, while NGD with CG is slower due to the many CG steps needed to approximate the Gramian inverse. BFGS is not competitive in iteration count, though its lower per-iteration cost allows for faster initial progress in wall-clock time. However, it ultimately stagnates at a higher error.
Figure 2: Poisson equation in 3D discretized via PINNs. From left to right: relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $500$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
5.2 Heat equation in 3+1D

### v3

@@ 4.2 NyströmNGD
We now present our method, NyströmNGD. To compute the regularized natural gradient, we approximately solve $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})\mathbf{d}=\nabla L(\boldsymbol{\theta})$ using the preconditioned conjugate gradient (PCG) method in a matrix-free fashion, with the preconditioner constructed via randomized Nyström. This approach leverages the structure of the problem: matrix-vector products with the Gramian $\mathbf{G}(\boldsymbol{\theta})$ can be computed efficiently via automatic differentiation, while the strong spectral decay commonly observed in practice (see Figure˜1) permits an accurate low-rank approximation, yielding a preconditioner that is both efficient to apply and effective in accelerating convergence. Additionally, the Nyström approximation provides estimates of the top $\ell$ eigenvalues of $\mathbf{G}$, which can be used to adapt the regularization parameter $\mu$ and the rank $\ell$ at each iteration. This section presents the resulting algorithm in full detail.
Adaptation of the regularization parameter $\mu$
@@
The regularization parameter $\mu>0$ is crucial for numerical stability and for mitigating the slow initial convergence typical of quasi-second-order methods like NGD. Prior works proposed heuristics that adapt $\mu$ dynamically [13, 33] or decrease it with the loss [85], so that early iterations mimic gradient descent and gradually shift to NGD near a local minimum. We instead propose a heuristic based on numerical rank of $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})$.
@@
Let $\lambda_{1}\geq\dots\geq\lambda_{p}$ denote the eigenvalues of the Gramian $\mathbf{G}(\boldsymbol{\theta})$. We set the regularization parameter as $\mu=\gamma\cdot\epsilon_{\mathrm{mach}}\cdot\lambda_{1}$, with $\gamma=p$ as a reasonable default. This is the threshold commonly employed to compute the numerical rank, i.e., the cut-off under which eigenvalues are considered to be numerically zero [45, 69]. Observe that although $\lambda_{1}$ is not directly available, it can be estimated by $\hat{\lambda}_{1}$ from randomized Nyström. When $\mathbf{G}(\boldsymbol{\theta})$ exhibits strong spectral decay, the expected relative error of this estimate is small even for modest $\ell$ [60, Theorem 14.1]. Finally, we have observed that it can be beneficial to enforce stronger regularization in early iterations, when still far from local minima. This can be achieved with a heuristic of the form $\mu=\max(\gamma\cdot\epsilon_{\mathrm{mach}}\cdot\hat{\lambda}_{1},c\cdot\left\|{\nabla L(\boldsymbol{\theta})}\right\|^{\alpha})$ or $\mu=\max(\gamma\cdot\epsilon_{\mathrm{mach}}\cdot\hat{\lambda}_{1},c\cdot|L(\boldsymbol{\theta})|^{\alpha})$, where the scaling $c$ and the exponent $\alpha$ depend on the problem.
Adaptation of rank parameter $\ell$
@@
The rank parameter $\ell$ controls the computational cost of the preconditioner and should ideally scale with the effective dimension $d_{\mathrm{eff}}(\mu)$ [31, Theorem 5.1]. As $d_{\mathrm{eff}}(\mu)$ is not directly accessible, we estimate $\ell$ adaptively. Following [31, Section 5.4.2], we do it by monitoring the ratio $\hat{\lambda}_{\ell}/\mu$: if it goes above $10$, we double $\ell$ for the next iteration, up to a prescribed maximum $\ell_{\max}$. Otherwise, we reduce $\ell$ to the smallest index satisfying $\hat{\lambda}_{\ell}/\mu>10$, adding a small offset (e.g., $1$ or $2$) to avoid oscillations.
@@
Algorithm˜2 summarizes the proposed algorithm. At each iteration, we construct the randomized Nyström preconditioner (lines 5 to 7) using the matrix-free Gramian. We then apply preconditioned conjugate gradient (PCG) to solve the linear system (6) (line 9). The PCG stopping criterion combines a maximum number of iterations maxit and a relative tolerance $\mathtt{rel\_tol}=\min(\kappa,\left\|{\nabla L(\boldsymbol{\theta}_{k})}\right\|_{2})$, inspired by the Steihaug–Toint truncated CG method [20]. This design avoids overspending computation far from critical points, where NGD steps may underperform compared to GD, and ensures accuracy near local minima. Once the search direction $\mathbf{d}_{k}$ is computed, a line search is performed to update the iterate (lines 10 and 11). In practice, we found that a standard backtracking line-search [83] suffices, without requiring logarithmic grid searches as in [64, 46]. Finally, the rank parameter $\ell$ is updated adaptively (lines 12 to 16).
@@
5:  $\mathbf{U},\hat{\boldsymbol{\Lambda}}]=\mathtt{RandomizedNystrom}(\mathbf{G}_{\mathrm{fun}},\ell) \triangleright$ Randomized Nyström approximation
6:  $\mu=\gamma\cdot\hat{\lambda}_{1}\cdot\epsilon_{\mathrm{mach}}$ $\triangleright$ Regularization parameter $\mu$
7:  $\mathbf{P}_{\mathrm{fun}}=\mathbf{v}\mapsto\frac{1}{\hat{\lambda}_{\ell}+\mu}\mathbf{U}(\hat{\boldsymbol{\Lambda}}+\mu\mathbf{I})\mathbf{U}^{\top}\mathbf{v}+(\mathbf{I}-\mathbf{U}\mathbf{U}^{\top})\mathbf{v}$ $\triangleright$ Matrix-free preconditioner
8:  $\mathtt{rel\_tol}=\min(\kappa,\left\|{\nabla L(\boldsymbol{\theta}_{k})}\right\|_{2})$ $\triangleright$ Relative tolerance for PCG
9:  $\mathbf{d}_{k}=\texttt{PCG}(\mathbf{G}_{\mathrm{fun}}+\mu\cdot\mathrm{Id},\,\nabla L(\boldsymbol{\theta}_{k}),\,\mathtt{rel\_tol},\,\mathtt{maxit},\,\mathbf{P}_{\mathrm{fun}})$ $\triangleright$ Run PCG
10:  $\alpha_{k}=\texttt{Linesearch}(\boldsymbol{\theta}_{k},-\mathbf{d}_{k})$ $\triangleright$ Linesearch along $\alpha\mapsto\boldsymbol{\theta}_{k}-\alpha\mathbf{d}_{k}$
@@
12:  if $\hat{\lambda}_{\ell}>10\mu$ then
13:    $\ell=\min(2\ell,\ell_{\max})$ $\triangleright$ Possibly increase rank parameter $\ell$
@@
15:    $\ell=\min\left(\min\{i:\hat{\lambda}_{i}<10\mu\}+1,\,\ell_{\max}\right)$ $\triangleright$ Possibly decrease rank parameter $\ell$
@@
Remark 4.4.
One key advantage of NyströmNGD over standard NGD lies in its reduced and tunable memory requirements. The most memory-intensive step is the Randomized Nyström approximation (line 5, using Algorithm˜1), which at most requires storing two $p\times\ell$ matrices in memory simultaneously. Consequently, the peak memory footprint corresponds to roughly $2\ell$ vectors of length $p$, comparable to the memory complexity of L-BFGS with a history size of $\ell$. This provides explicit control over memory consumption through the parameter $\ell_{\max}$, enabling the method to scale to larger networks. Naturally, smaller values of $\ell_{\max}$ may yield a weaker preconditioner and increase the number of CG iterations, yet the algorithm remains fully operational.
Remark 4.5.
The Randomized Nyström method is not the only viable approach for constructing a symmetric positive definite preconditioner in a matrix-free manner. Alternatives include techniques based on partial Cholesky factorizations. Since the Gramian $\mathbf{G}$ is dense, Incomplete Cholesky is unsuitable; however, one could employ Pivoted Partial Cholesky, either in its deterministic form [43] or its randomized variant [17]. The main limitation of both approaches is that they require access to the full diagonal of $\mathbf{G}$: the deterministic greedy pivoting in [43] and the randomized RPCholesky method in [17] both rely on it. In our context, this involves computing matvecs of $\mathbf{G}$ with all canonical basis vectors. This has the same FLOP cost as constructing the full $\mathbf{G}$, though a lower memory usage. For this reason, we do not pursue these methods further and only include a brief numerical comparison in D.
@@ 5 Numerical Experiments
We test NyströmNGD on a range of PDE problems with known analytical solutions. Both strong formulations, discretized with PINNs, and weak formulations, discretized using FEINNs, are considered. Additional numerical experiments are provided in appendix: a sensitivity analysis of NyströmNGD with respect to its key hyperparameters is provided in C, and a brief comparison with alternative preconditioners to randomized Nyström is presented in D.
@@
We compare NyströmNGD against Natural Gradient Descent with an SVD-based pseudoinverse (legend: NGD), matrix-free NGD solved via conjugate gradient without preconditioning (legend: NGD with CG), and the BFGS algorithm [83]. NGD-based methods are run for $500$ or $1000$ iterations, depending on the example, while BFGS is run for $6\times 10^{4}$ or $1.2\times 10^{5}$ iterations. For a fair comparison with NyströmNGD, NGD with CG uses the same relative tolerance as in Algorithm˜2, with a maximum number of iterations set to $\texttt{maxit}+\ell_{\max}$. Given the differing per-iteration costs of the methods, we compare performance in terms of both iteration count and wall-clock time. First-order optimizers (e.g., gradient descent, Adam) and L-BFGS are excluded, as they have been shown to be non-competitive; see Tables˜2 and 1 and [46, 65]. Unless otherwise stated, all experiments use fully connected neural networks with $\tanh$ activation and two hidden layers of width $32$. Each experiment is averaged over 10 runs, with solid lines in the plots indicating medians and shaded regions representing the first and third quantiles.
5.1 PDEs in strong form: PINNs
We begin with PDEs in strong form discretized via Physics-Informed Neural Networks (PINNs) [70, 21].
@@ 5.1 PDEs in strong form: PINNs
In all PINN experiments, we set in Algorithm˜2 the parameters $\ell_{0}=10$, $\ell_{\max}=500$, $\gamma=p$, $\mathtt{maxit}=20$, and $\kappa=0.1$. For NGD and NGD with CG, we use $\mu=\min(10^{-5},L(\boldsymbol{\theta}))$ as suggested in [85], from which most experiments are adapted. To enforce stronger regularization in early iterations, NyströmNGD employs $\mu=\max(p\cdot\hat{\lambda}_{1}\cdot\epsilon_{\mathrm{mach}},10^{-4}\cdot L(\boldsymbol{\theta})^{2})$. Unless stated otherwise, integrals in the loss and metric are discretized via Monte Carlo quadrature with $N_{\Omega}=3000$ points uniformly sampled in the interior of the domain, and $N_{\Gamma}=800$ points on the boundary. Random integration points are sampled once and fixed throughout training. Errors are always computed using an order of magnitude more quadrature points than during training, sampled independently.
For the PINN formulations, we use JAX [16] and JAXopt [12]. The JAXopt BFGS implementation was optimized for efficiency, as detailed in B, achieving, together with JIT compilation, a $1000\times$ speedup over the timings reported in [64]. The NGD implementation is taken from the public repository of [85]. Integrals in both the loss and the metric are discretized via Monte Carlo quadrature. All experiments are conducted in double precision on a single NVIDIA RTX 4090 GPU with 24 GB of memory.
A comparison with other optimizers
As already noted, first-order optimizers and L-BFGS have been shown to be non-competitive compared to NGD and BFGS [46, 65]. To experimentally confirm this, we include a brief comparison of NyströmNGD and BFGS with other state-of-the-art algorithms: L-BFGS [83], Adam [51], Shampoo [35, 4], and Muon [47]. For L-BFGS, Adam, and Muon, we use the Optax implementation [26]; for Shampoo, we use the implementation by Google Research. Results are reported in Tables˜1 and 2 for the three problems discussed later in detail.
Our proposed method consistently outperforms the other optimizers in terms of final accuracy, even when they are given a larger computational time budget. Moreover, as shown in Table˜2, NyströmNGD reaches the same accuracy as other methods in substantially less time. These results clearly indicate that the best runner-up to NyströmNGD is the BFGS method, which is therefore the only one included in the subsequent comparisons.
Table 1: PDEs in strong form discretized via PINNs. Final relative $\mathrm{H}^{1}$ error, for different optimizers.
Poisson Heat Kovasznay
NyströmNGD $2.92\text{\times}{10}^{-05}$ $2.34\text{\times}{10}^{-05}$ $3.01\text{\times}{10}^{-06}$
BFGS $1.88\text{\times}{10}^{-04}$ $34\text{\times}{10}^{-04} 1.1$$10\text{\times}{10}^{-05}$
L-BFGS $8.28\text{\times}{10}^{-04}$ $4.42\text{\times}{10}^{-04}$ $7.93\text{\times}{10}^{-05}$
Adam $2.19\text{\times}{10}^{-02}$ $6.77\text{\times}{10}^{-03}$ $4.18\text{\times}{10}^{-03}$
Shampoo $8.44\text{\times}{10}^{-02}$ $2.52\text{\times}{10}^{-02}$ $8.84\text{\times}{10}^{-01}$
Muon $2.84\text{\times}{10}^{-02}$ $7.78\text{\times}{10}^{-03}$ $1.48\text{\times}{10}^{-03}$
Table 2: PDEs in strong form discretized via PINNs. Total training time (in seconds) for different optimizers; NyströmNGD’s time to reach the same accuracy is shown in parentheses.
Poisson Heat Kovasznay
NyströmNGD 86.1 335.4 485.2
BFGS 81.7 (11.1) 304.6 (22.0) 363.6 (226.5)
L-BFGS 295.4 (9.7) 685.7 (19.4) 476.3 (189.0)
Adam 228.5 (5.3) 911.4 (13.7) 504.7 (97.9)
Shampoo 503.5 (1.9) 1052.0 (6.9) 766.6 (0.8)
Muon 277.5 (4.6) 951.9 (12.7) 635.2 (127.3)
5.1.1 Poisson problem in 3D
@@
with exact solution $u(x,y,z)=\sin(\pi x)\sin(\pi y)\sin(\pi z)$, as in [85]. Problem (13) is linear and is considered in strong form. Using the notation of Section˜2.2.1, the function spaces are $\mathrm{V}=\mathrm{H}^{2}(\Omega)$, $\mathrm{X}=\mathrm{L}^{2}(\Omega)$, and $\mathrm{Y}=\mathrm{L}^{2}(\partial\Omega)$, with operators
$\mathcal{F}(u)=-\Delta u,\qquad\mathcal{B}(u)=\left.{u}\right|_{\partial\Omega}.$
The corresponding least-squares loss is
and the induced metric reads
Remark 5.1.
The metric (15) is coercive only in the $\mathrm{H}^{1/2}(\Omega)$ norm, but not in $\mathrm{H}^{2}(\Omega)$. This is due to the use of the $\mathrm{L}^{2}(\partial\Omega)$ norm for the Dirichlet boundary term. Coercivity in $\mathrm{H}^{2}(\Omega)$ can be restored by enforcing the boundary conditions strongly [85, Remark 6] or by penalizing in the $\mathrm{H}^{1/2}(\partial\Omega)$ norm [14]. However, in line with standard PINN practice, we retain the $\mathrm{L}^{2}(\partial\Omega)$ penalty for simplicity.
Results
@@
As shown in Figure˜2, NyströmNGD yields the best results in both final accuracy and computational efficiency, converging in around $100$ iterations. The Nyström rank parameter $\ell$ increases rapidly during early iterations, consistent with Figure˜3, and then stabilizes well below the maximum allowed rank $\ell_{\max}=500$. This confirms the effectiveness of the adaptive rank selection.
NGD initially converges faster per iteration but quickly plateaus. According to [85], it would require roughly $5000$ iterations to reach the accuracy achieved by NyströmNGD, resulting in a substantially higher total cost. Importantly, NyströmNGD not only lowers computational cost but also improves final accuracy compared to standard NGD. While we do not have a definitive explaination, this behavior may stem from enhanced numerical stability or from the stochastic nature of the Nyström approximation, which can help escape poor local minima, analogous to how SGD can outperform full-batch GD in classical machine learning [53, 48, 40]. NGD with CG performs similarly to NGD in iteration count but is considerably slower due to the large number of CG steps needed to approximate the Gramian inverse, underscoring the crucial role of preconditioning. BFGS is not competitive in iteration count; although its low per-iteration cost yields faster initial progress in wall-clock time, it ultimately stagnates at a higher error.
(a) Relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $500$ iterations)
(b) Relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time.
(c) Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
Figure 2: Convergence for the 3D Poisson equation discretized via PINNs.
Figure 3: 3D Poisson equation discretized via PINNs and trained with NGD using an SVD-based pseudoinverse. Left: decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. Right: relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations.
5.1.2 Scalability
Although an MLP with two hidden layers of 32 neurons each was already sufficient to achieve a relative accuracy on the order of $10^{-5}$, limited primarily by quadrature error, we also assess the scalability of NyströmNGD. As discussed in Remark˜4.4, NyströmNGD provides explicit memory control via the parameter $\ell_{\max}$, enabling its application to significantly larger networks than standard NGD.
We perform both width ($32\to 256$) and depth ($2\to 6$) sweeps for the 3D Poisson problem. We set $\ell_{\max}=500$, except for the networks with $128$ and $256$ neurons per layer, where we use $\ell_{\max}=300$ and $\ell_{\max}=150$, respectively. We compare NyströmNGD against BFGS and L-BFGS. Results are summarized in Tables˜3, 4, 5 and 6.
NyströmNGD scales effectively to networks substantially larger than those trainable with standard NGD, consistently outperforming both BFGS and L-BFGS in terms of final accuracy ( Tables˜3 and 5) and total runtime ( Tables˜4 and 6). As expected, decreasing $\ell_{\max}$ leads to a weaker preconditioner and an increased number of CG iterations, as shown in Tables˜7 and 8, yet NyströmNGD continues to provide the best trade-off between computational cost and accuracy. For instance, on the largest tested network (width 256, 67k params), BFGS runs out of memory on our GPU. In contrast, NyströmNGD, with memory usage comparable to L-BFGS, reaches a lower error with a $7\times$ speedup for the same accuracy. These findings underscore the improved scalability of NyströmNGD over standard NGD.
Table 3: Width sweep ($32\to 256$) on the 3D Poisson problem discretized via PINNs. Final relative $\mathrm{H}^{1}$ errors.
$\boldsymbol{32\times 2}$ (1.2k) $\boldsymbol{64\times 2}$ (4.5k) $\boldsymbol{128\times 2}$ (17k) $\boldsymbol{256\times 2}$ (67k)
NyströmNGD $2.37\text{\times}{10}^{-05}$ $1.28\text{\times}{10}^{-05}$ $1.83\text{\times}{10}^{-05}$ $7.53\text{\times}{10}^{-05}$
BFGS $1.86\text{\times}{10}^{-04}$ $7.07\text{\times}{10}^{-05}$ $5.59\text{\times}{10}^{-05}$ Out Of Memory
L-BFGS $1.32\text{\times}{10}^{-03}$ $7.12\text{\times}{10}^{-04}$ $6.00\text{\times}{10}^{-04}$ $4.57\text{\times}{10}^{-04}$
Table 4: Width sweep ($32\to 256$) on the 3D Poisson problem discretized via PINNs. Total optimization times (in seconds); NyströmNGD’s time to reach the same accuracy is shown in parentheses.
$\boldsymbol{32\times 2}$ (1.2k) $\boldsymbol{64\times 2}$ (4.5k) $\boldsymbol{128\times 2}$ (17k) $\boldsymbol{256\times 2}$ (67k)
NyströmNGD 38.2 94.6 247.1 541.6
BFGS 71.4 (15.4) 125.3 (20.6) 678.5 (51.2) Out Of Memory
L-BFGS 204.8 (11.2) 260.9 (15.7) 378.0 (34.4) 762.5 (111.8)
Table 5: Depth sweep ($2\to 6$) on the 3D Poisson problem discretized via PINNs. Final relative $\mathrm{H}^{1}$ errors.
$\boldsymbol{32\times 2}$ (1.2k) $\boldsymbol{32\times 3}$ (2.3k) $\boldsymbol{32\times 4}$ (3.3k) $\boldsymbol{32\times 5}$ (4.4k) $\boldsymbol{32\times 6}$ (5.4k)
NyströmNGD $37\text{\times}{10}^{-05} 2.2$$21\text{\times}{10}^{-05}$ $2.59\text{\times}{10}^{-05}$ $4.35\text{\times}{10}^{-05}$ $2.10\text{\times}{10}^{-05}$
BFGS $1.86\text{\times}{10}^{-04}$ $1.42\text{\times}{10}^{-04}$ $1.25\text{\times}{10}^{-04}$ $9.52\text{\times}{10}^{-05}$ $8.70\text{\times}{10}^{-05}$
L-BFGS $32\text{\times}{10}^{-03} 1.1$$17\text{\times}{10}^{-03}$ $1.21\text{\times}{10}^{-03}$ $50\text{\times}{10}^{-03} 1.1$$18\text{\times}{10}^{-03}$
Table 6: Depth sweep ($2\to 6$) on the 3D Poisson problem discretized via PINNs. Total optimization times (in seconds); NyströmNGD’s time to reach the same accuracy is shown in parentheses.
$\boldsymbol{32\times 2}$ (1.2k) $\boldsymbol{32\times 3}$ (2.3k) $\boldsymbol{32\times 4}$ (3.3k) $\boldsymbol{32\times 5}$ (4.4k) $\boldsymbol{32\times 6}$ (5.4k)
NyströmNGD 38.2 59.4 80.6 87.5 121.7
BFGS 71.4 (15.4) 107.5 (13.3) 153.5 (22.8) 199.8 (29.5) 247.4 (24.6)
L-BFGS 204.8 (11.2) 248.5 (10.6) 291.2 (17.0) 341.1 (23.5) 382.1 (17.2)
Table 7: Width sweep ($32\to 256$) on the 3D Poisson problem discretized via PINNs. Average Nyström rank parameter $\ell$ (with $\ell_{\max}$ in parentheses) and number of CG iterations for NyströmNGD.
$\boldsymbol{32\times 2}$ (1.2k) $\boldsymbol{64\times 2}$ (4.5k) $\boldsymbol{128\times 2}$ (17k) $\boldsymbol{256\times 2}$ (67k)
[…]
CG it. 13 17 22 38
Table 8: Depth sweep ($2\to 6$) on the 3D Poisson problem discretized via PINNs. Average Nyström rank parameter $\ell$ (with $\ell_{\max}$ in parentheses) and number of CG iterations for NyströmNGD.
$\boldsymbol{32\times 2}$ (1.2k) $\boldsymbol{32\times 3}$ (2.3k) $\boldsymbol{32\times 4}$ (3.3k) $\boldsymbol{32\times 5}$ (4.4k) $\boldsymbol{32\times 6}$ (5.4k)
[…]
CG it. 13 15 16 16 17
5.1.3 Heat equation in 3+1D